Flamo:

Solve the system y = -x + 7 and y = 0.5(x - 3)2

Hero:

@Flamo multiply both sides of y = -x + 7 by 2.

Flamo:

2y = -2x + 14

Hero:

Next, multiply both sides of \(y = 0.5(x - 3)^2\) by 2.

Flamo:

2y = 1(2x - 6)^2?

Hero:

More like \(2y = 1(x - 3)^2\)

Flamo:

Oh.

Hero:

Next you set both expressions for 2y equal to each other.

Flamo:

Hmm..

Flamo:

I have this written down. y = -x + 7 and y = 0.5(x - 3)^2 0.5(x - 3)^2 = -x + 7 Multiply Both sides by 2 -2x + 14 = x^2 - 6x + 9

Flamo:

So that's incorrect?

Hero:

It's the exact same thing actually.

Flamo:

Oh, So what do you need to do *from there*?

Hero:

Solve for x, which is what we were doing previously.

Flamo:

Oh

Hero:

I was trying to show you how to do it but once again, you have introduced something else which is adding to your confusion.

Flamo:

I did? Oh...

Flamo:

The Teacher says, After -2x + 14 = x^2 - 6x + 9 <---- You need to Simplify: x^2 - 4x - 5 = 0 then factor to figure out the x's, then sub it x's to get y

Hero:

Yeah, see you did it again. I was showing you how to do it, then you interrupt by introducing "what the teacher said". I don't care what the teacher said.

Flamo:

Oh, You said Solve for x. Right?

Flamo:

No no no no, You said: Next you set both expressions for 2y equal to each other.

Flamo:

Right?

Hero:

Yes, I did and then you proceeded to post what your teacher said.

Flamo:

Oh, Sorry. I was Confused. Let's Continue from there then.

Hero:

Okay, still waiting for you to set the expressions for 2y equal to each other.

Flamo:

Yeah, i'm doing that right now.

Flamo:

2y = -2x + 14 and 2y = 1(x - 3)^2

Flamo:

Right?

Hero:

Yes, and so since we know 2y = 2y, then what expressions are equal to each other?

Flamo:

Uhh, I don't know.. ;-;

Hero:

\(2y = 2y\) means that \(-2x + 14 = (x - 3)^2\)

Flamo:

OH

Hero:

2y = -2x + 14 || || 2y = 1(x - 3)^2

Flamo:

So You Subtract?

Hero:

"Wondering how @flamo concluded that equal signs mean subtract".

Flamo:

Oh... They were Sideways, Didn't realize that they were equal Signs...

Flamo:

My bad... ;-;

Flamo:

Okay, the equations are equal, now what? ;-;

Hero:

You have to post your steps here.

Flamo:

y = -x + 7 and y = 0.5(x - 3)^2 0.5(x - 3)^2 = -x + 7 Multiply Both sides by 2 2y = -2x + 14 and 2y = 1(x - 3)^2 2y = -2x + 14 = 2y = 1(x - 3)^2

Flamo:

That's the steps I've written when we were solving the problem

Hero:

It's true what you've written except since we need to solve for \(x\) you only include the expressions with \(x\)

Flamo:

Oh? So You take the 2y out on Both equations?

Hero:

Yes

Flamo:

-2x + 14 = 1(x - 3)^2? Sorry I had to do something, ;-;

Flamo:

@Hero

Hero:

Finally.... Now question. what is \(1 \times 7 = \)

Flamo:

7

Hero:

Good so \(1 \times (x - 3)^2\) =

Flamo:

...

Flamo:

Oh.

Flamo:

Uhh (x - 3)^2

Hero:

Correct. So basically what we have is \(-2x + 14 = (x - 3)^2\)

Hero:

And we can expand \((x - 3)^2\) using (not FOIL) but distributive property

Flamo:

Oh

Flamo:

So (x - 3) - (3 - x)?

Hero:

No \((x - 3)^2\) means multiply \((x - 3)\) by itself twice.

Flamo:

Oh... I feel Dumb. ;-; \[(x - 3) \times (x - 3)\]

Flamo:

Which is \[x^2 - 6x +9 \] Right @Hero?

Hero:

Correct

Flamo:

y = -x + 7 and y = 0.5(x - 3)^2 0.5(x - 3)^2 = -x + 7 Multiply Both sides by 2 2y = -2x + 14 and 2y = 1(x - 3)^2 -2x + 14 = 1(x - 3)^2 -2x + 14 = (x - 3)^2 -2x + 14 = (x - 3) * (x - 3) -2 + 14 = x^2 - 6x + 9

Flamo:

This is what I have so Far. Correct?

Flamo:

@Hero

Hero:

Looks good

Flamo:

Okay, So now what?

Flamo:

From this: -2 + 14 = x^2 - 6x + 9

Hero:

Actually the x is missing from 2.

Flamo:

Oops

Flamo:

From this: -2x + 14 = x^2 - 6x + 9

Hero:

Next subtract 14 from both sides. Then add 2x to both sides.

Flamo:

\[-2x = x^2 - 6x + (-5)\]

Flamo:

Okay, I subtracted 14 from both Sides, now add 2x?

Flamo:

Or Subtract *-2x*?

Flamo:

From Both sides.

Hero:

It's exactly like I suggested above.

Flamo:

Oh, So \[x^2 - 4x + (-5)\]

Flamo:

Okay, So I added 2x to both sides.

Hero:

Yeah but you still need to put something on the other side of the equal sign.

Flamo:

0?

Flamo:

0 = x^2 - 4x + (-5)?

Hero:

Yes correct

Flamo:

Oh, Okay. Now what?

Hero:

You can re-write the expression above as \(0 = x^2 - 4x - 5\) since \(a + (-b) = a - b\)

Flamo:

Oh, Okay

Flamo:

Okay, I did that, Now what?

Flamo:

\[0 = x^2 - 4x + 5\]

Hero:

You wrote the equation I wrote above incorrectly.

Flamo:

Oh.. ;-;

Flamo:

\[0 = x^2 - 4x - 5\]

Flamo:

Okay, now what?

Hero:

Now you factor \(x^2 - 4x - 5\)

Flamo:

(x - 5) (x + 1)

Hero:

Yeah, but what are your steps to get that?

Flamo:

AC Method?

Hero:

When I ask "What are your steps?", that implies for you to actually post those steps.

Flamo:

Oh, I used a Calculator, ;-;

Flamo:

Oh, I used a Calculator. ;-;

Hero:

Okay, I'm about to go take a nap.

Flamo:

...

Flamo:

;-;

Flamo:

I can't stop you... ;/

Hero:

You have to learn how to factor quadratic expressions. It takes too much time to teach. Review some of the other threads where I've shown other users how. I know you've been watching. Don't pretend you haven't.

Flamo:

Yes, I have been watching.

Hero:

We have to find two numbers \(m\) and \(n\) such that \(m + n = |b|\) \(m \times n = ac\) We need to find that so that we can input \(m + n\) here in place of \(b\) \(x^2 -(m + n)x - 5\) Obviously \(m + n = 4\) in this case.

Flamo:

\[x^2 - (4)x - 5\]

Flamo:

Or Just 4x

Hero:

Yes we know that already, but it doesn't get us any closer in terms of factoring. We have to "split the middle term"

Flamo:

2 + 2?

Flamo:

= 4?

Flamo:

2 + 2 are the middle Terms for M and n

Hero:

No, \(m \times n\) has to equal -5

Hero:

\(m + n\) has to equal \(4\)

Flamo:

Oh

Flamo:

(m + n) = 4 So m = 2 as well as n. So (2+2) = 4.

Hero:

\(m\) and \(n\) has to satisfy both requirements. How about \(5 - 1 = 4\) \((5)(-1) =-5\)

Hero:

So looks like it's going to be \(m = 5\) and \(n = -1\)

Flamo:

Oh.

Flamo:

\[x^2 - 5 + (-1) - 5\]?

Flamo:

Or do You not put it in the equation?

Hero:

You forgot the x. It's \(x^2 -(5 - 1)x - 5\)

Flamo:

Oh..

Hero:

And the whole time, the expression should be set equal to zero.

Flamo:

Oh

Flamo:

\[0 = x^2 - (5 - 1)x - 5\]?

Hero:

Yes, now you have to distribute the negative and the x within the parentheses.

Flamo:

Oh.. Hm..

Flamo:

I don't get it.. Sorry ;-;

Hero:

Well, I'm going to take a nap. Review the tutorials.

Flamo:

Okay. Have fun... ;-;

Flamo:

I suck.. ;/

Flamo:

...

Flamo:

x^2 - 4x - 5

Flamo:

x^2 = x * x

Flamo:

(x )(x )

Flamo:

-5 = - 5

Flamo:

(x - 5)(x - 1)

Flamo:

):

Flamo:

(x - 5)(x + 1)*

Flamo:

;-;

Flamo:

@Hero you still here?

Flamo:

Never Mind, I took enough of your time anyway

Flamo:

(x - 5)(x + 1)

Flamo:

Check

Flamo:

Hm..

Flamo:

x(x - 1) + 5(x - 1)

Flamo:

x^2 - x + 5x - 5

Flamo:

-x subtract on 5x

Flamo:

x^2 -4x - 5

Flamo:

Oh.

Flamo:

@Shadow